# fMoE: Fine-Grained Expert Offloading for Large Mixture-of-Experts Serving

### v1

@@ 4.4 Expert Map Store Management
Theoretical analysis. The expert map deduplication can be formulated as a Minimum Sphere Covering problem [17]. The objective is to minimize the total number of expert maps in the store, where each expert map is a vector representation of spheres, while maximizing the sphere coverage in the expert activation space. Studies [41, 15] have proved that maintaining at least expert maps guarantees a lower bound of 75% expert map similarity (i.e., we can find an expert map that is at least 75% similar to any new iterations), and keeping expert maps provides a lower bound of 98% similarity, where and are the numbers of layers and experts per layer in the MoE model, respectively. Given that modern MoE-based LLMs generally have and , we can approximate the Expert Map Store’s maximal requirement to be less than 50K expert maps with 200 MB CPU memory [54].
4.5 Expert Caching and Eviction
Similar to existing expert offloading solutions [16, 54, 47], we design fMoE to maintain an Expert Cache on GPUs to reuse expert weights when serving different request prompts. Given matched expert maps from §4.2, we guide fMoE’s Expert Cache to compute two priority scores for individual experts: 1) a prefetching priority to decide the prefetching orders of experts in the searched maps, and 2) an eviction priority to determine the eviction orders of experts in the Expert Cache.
@@ 4.5 Expert Caching and Eviction
where is the expert probability from the searched expert map, and is the current layer that inference process stays at. Intuitively, experts with higher probability to be activated should be prefetched sooner, and experts that sit closer to the current layer (i.e., smaller ) should also be prioritized.
@@
Expert eviction priority. Similar to MoE-Infinity [54], fMoE’s expert caching is based on the least frequently used (LFU) algorithm. We integrate the searched map to jointly determine the eviction priority. For each expert , we define the eviction priority to be
@@
where is the cache visit frequency and is the probability from the searched map for an expert . Intuitively, when reaching the Expert Cache limit, we want to first evict experts who are less frequently hit and have lower probabilities of being activated. Note that similar to existing works [54, 47], we do not consider the recent usage of experts as opposed to the classic least recently used (LRU) algorithm [16]. Since the expert usage is layer-wise sequential, i.e., one layer following another, prioritizing recently used experts is against the nature of sequential forward computation in MoE serving.
@@
On-demand expert loading. Mispredictions of expert prefetching lead to expert miss in the Expert Cache, as the MoE model cannot find available experts designated by the gate networks. Whenever an expert miss occurs, fMoE pauses all expert prefetching tasks and immediately loads missed experts from CPU to GPU memory for fast serving.
5 fMoE’s Implementation
@@
We prototype fMoE on top of Huggingface Transformers framework [52] using MoE-Infinity codebase [55]. The implementation of fMoE is described as follows.
@@
Expert Map Store is implemented in Python using PyTorch [37] and NumPy [19] libraries. We store all semantic embeddings and expert maps using ndarrays data structure for efficient array operations. The arrays are converted to tensors to compute similarity for expert map matching.
Expert Map Matcher is implemented in Python using PyTorch [37] and TorchMetrics [12] libraries. We implement the pairwise computations, including similarity (§4.2) and redundancy (§ 4.4) scores, using the Cosine Similarity interfaces in TorchMetrics. We use the Python multithreading library to implement the asynchronous expert map matching and expert prefetching, where the threads share the same memory space with the Expert Map Store for efficient reading and writing.
@@
Expert Cache is implemented in C++ based on MoE-Infinity codebase [55]. The expert management in GPUs is implemented with the CUDA Runtime APIs [35]. We implement the caching logic of fMoE and fix critical bugs in the MoE-Infinity codebase to enable expert offloading. Same with MoE-Infinity, fMoE supports multi-GPU inference with expert parallelism, where the experts are mapped to different GPU devices for loading and offloading. We use a hash map to assign expert IDs to different GPUs and retrieve them during inference. The expert assignment follows a round-robin manner to balance the overall GPU load. Additionally, we use a multi-thread task pool in the GPU space to schedule and execute expert prefetching and on-demand loading tasks.
6 Evaluation
MoE Models | Parameters | Experts Per Layer | Num. of |
(active / total) | (active / total) | Layers | |
Mixtral-87B [21] | 12.9B / 46.7B | 2 / 8 | 32 |
Qwen1.5-MoE [57] | 2.7B / 14.3B | 4 / 60 | 24 |
Phi-3.5-MoE [1] | 6.6B / 42B | 2 / 16 | 32 |
6.1 Experimental Setup
@@
Testbed. We conduct all experiments on a six-GPU testbed, where each GPU is an NVIDIA GeForce RTX 3090 with 24 GB GPU memory. All GPUs are inter-connected using pairwise NVLinks and connected to the CPU memory using PCIe 4.0 with 32GB/s bandwidth. Additionally, the testbed has a total of 32 AMD Ryzen Threadripper PRO 3955WX CPU cores and 480 GB CPU memory.
@@
Models. We employ three popular MoE-based LLMs in our evaluation: Mixtral-87B [21], Qwen1.5-MoE [57], and Phi-3.5-MoE [1]. Table 1 describes the parameters, number of MoE layers, and number of experts per layer for the three models. Following the evaluation of existing works [47], we profile the models to set the optimal prefetch distance to three before evaluation.
@@
Datasets and traces. We employ two real-world prompt datasets commonly used for LLM evaluation: LMSYS-Chat-1M [60] and ShareGPT [45]. For most experiments, we split the sampled datasets in a standard 7:3 ratio, where 70% of the prompts’ context data (i.e., semantic embeddings and expert maps) are stored in fMoE’s Expert Map Store, and 30% of the prompts are used for testing. For online serving experiments, we empty the Expert Map Store and use real-world LLM inference traces [38, 48] released by Microsoft Azure to set input and generation lengths and drive invocations.
@@
Baselines. We compare fMoE against four SOTA MoE serving baselines: 1) MoE-Infinity [54] uses coarse-grained request-level expert activation patterns and synchronous expert prediction and prefetching for MoE serving. We prepare the expert activation matrix collection for MoE-Infinity before evaluation for a fair comparison. 2) ProMoE [47] employs a stride-based speculative expert prefetching approach for MoE serving. Since the codebase of ProMoE is not open-sourced and requires training predictors for each MoE model, we reproduced a prototype of ProMoE on top of MoE-Infinity in our best effort. 3) Mixtral-Offloading [16] combines a layer-wise speculative expert prefetching and a LRU-based expert cache. 4) DeepSpeend-Inference employs an expert-agnostic layer-wise parameter offloading approach, which uses pure on-demand loading and does not support prefetching. We implement the offloading logic of DeepSpeed-Inference in the MoE-Infinity codebase and add an expert cache for a fair comparison. We enable all baselines to serve MoE models from HuggingFace Transformer [52].
@@
Metrics. Following the standard evaluation methodology of existing works [47, 54, 61, 3] on LLM serving, we report the performance of the prefill and decode stages separately. We measure Time-to-First-Token (TTFT) for the prefill stage and Time-Per-Output-Token (TPOT) for the decode stage. Additionally, we also report other system metrics, such as expert hit rate and overheads, for detailed evaluation.
6.2 Overall Performance
We first evaluate the performance of prefill and decode stages when running fMoE and other baselines with the three MoE models, where we measure Time-To-First-Token (TTFT) and Time-Per-Output-Token (TPOT) for each stage. Note that the inference latency with expert offloading tends to be higher than no offloading due to two reasons: 1) During inference, an excessive amount of parameters in MoE models are loaded and offloaded, which prolongs the inference latency. 2) All baselines and fMoE are implemented on top of the MoE-Infinity codebase [55], whose inference latency is inherently impacted by MoE-Infinity’s implementation. Nevertheless, comparing fMoE and baselines is fair with the same experimental setup.
@@ 6.2 Overall Performance
Figure 9 shows the TTFT, TPOT, and expert hit rate of fMoE and other four baselines when serving three MoE models with LMSYS-Chat-1M and ShareGPT datasets, respectively. DeepSpeed has both the worst TTFT and TPOT due to expert-agnostic offloading and lacking expert prefetching. While Mixtral-Offloading, ProMoE, and MoE-Infinity perform better than DeepSpeed-Inference, they are underperformed by fMoE because of coarse-grained offloading designs. Compared to DeepSpeed-Inference, Mixtral-Offloading, ProMoE, and MoE-Infinity, our fMoE reduces the average TTFT by 44%, 35%, 33%, 30%, and reduces the average TPOT by 70%, 61%, 55%, 48%, across three MoE models. For expert hit rate, Mixtral-Offloading achieves a higher hit rate than the other three baselines because of its synchronous speculative prefetching with a prefetch distance of 1. However, due to synchronous prefetching, its TTFT and TPOT are worse than others except DeepSpeed-Inference. fMoE improves the average expert hit rate by 147%, 11%, 34%, and 63% over DeepSpeed-Inference, Mixtral-Offloading, ProMoE, and MoE-Infinity, respectively.
6.3 Online Serving Performance
@@
Except for the offline evaluation (i.e., Expert Map Store in full capacity before serving), we also evaluate fMoE against other baselines in online serving settings. We empty the Expert Map Store of fMoE and the expert activation matrix collection of MoE-Infinity for the online serving experiment. The request traces are derived from Azure LLM inference traces [38, 48], with 64 requests randomly sampled to drive LMSYS-Chat-1M prompts for each MoE model serving. To ensure consistency, fMoE and all baselines input and generate the exact number of tokens specified in the traces. Figure 10 illustrates the CDF of end-to-end request latency across three MoE models. The results demonstrate that fMoE significantly reduces overall request latency compared to other baselines in online serving scenarios.
6.4 Impact of Expert Cache Limits
@@
We measure the TPOT of fMoE and other baselines by limiting the expert cache memory budget to investigate their performance in the latency-memory trade-off (§2.3). We mainly focus on TPOT to show the end-to-end performance impacted by varying cache limits. Figure 11 shows the TPOT of fMoE and other four baselines when serving three MoE models under different expert cache limits. We gradually increase the GPU memory allocated for caching experts from 6 GB to 96 GB while employing the same experimental setting in §6.2. Similarly, DeepSpeed-Inference has the worst TPOT due to being expert-agnostic. fMoE consistently outperforms Mixtral-Offloading, ProMoE, and MoE-Infinity under varying expert cache limits. Especially for limited GPU memory sizes (e.g., 6GB), fMoE reduces the TPOT by 32%, 24%, 18%, and 18%, compared to DeepSpeed-Inference, Mixtral-Offloading, ProMoE, and MoE-Infinity, across three MoE models, respectively. With fine-grained expert offloading, fMoE significantly reduces the expert on-demand loading latency while maintaining a lower GPU memory footprint, therefore achieving a better spot in the latency-memory trade-off of MoE serving.
6.5 Ablation Study
We present the ablation study of fMoE’s design.
@@ 6.5 Ablation Study
Effectiveness of expert map search. One of fMoE’s key designs is the expert map, which tracks expert selection preferences in fine granularity. We evaluate the effectiveness of the expert map against five expert pattern-tracking approaches as follows. 1) Speculate: speculative prediction used by Mixtral-Offloading [16] and ProMoE [47], 2) Hit count: request-level expert hit count used by MoE-Infinity [54], 3) Map (T): expert map with only trajectory similarity search, 4) Map (T+S): expert map with both trajectory and semantic similarity search, and 5) Map (T+S+): expert map with full features enabled, including trajectory and semantic similarity search (§4.2) and dynamic expert selection (§4.3). We implement the above methods in fMoE’s Expert Map Matcher for a fair comparison. Figure 12(a) shows the expert hit rate of the above expert pattern tracking methods. Speculative prediction is effective due to the widespread presence of residual connections in Transformer blocks. However, its effectiveness decreases drastically as prefetch distance increases [47]. The request-level expert activation count has the worst performance due to coarse granularity. As features are incrementally restored to fMoE’s expert map, the expert hit rate gradually increases, demonstrating its effectiveness.
@@
Effectiveness of expert prefetching and caching. We evaluate fMoE’s expert prefetching and caching against two caching algorithms: 1) LRU used by Mixtral-Offloading [16] and 2) LFU used by MoE-Infinity [54]. Figure 12(b) depicts the expert hit rate of fMoE and two baselines. The results show that LRU performs poorly in expert offloading scenarios. Though LFU achieves a higher hit rate than LRU, fMoE surpasses both, achieving the highest expert hit rate.
6.6 Sensitivity Analysis

### v2

@@ 4.4. Expert Map Store Management
Theoretical analysis. The expert map deduplication can be formulated as a Minimum Sphere Covering problem (Elzinga and Hearn, 1972). Each expert map is a vectorized patch, and the full sphere represents all possible expert selections. The objective is to cover as much of the sphere as possible using a small number of maps, keeping storage overhead low. Studies (Rankin, 1947; Dumer, 2007) have proved that maintaining at least expert maps guarantees a lower bound of 75% expert map similarity (i.e., we can find an expert map that is at least 75% similar to any new iterations), and keeping expert maps provides a lower bound of 98% similarity, where and are the numbers of layers and experts per layer in the MoE model, respectively. Given that modern MoE-based LLMs generally have and , we can approximate the Expert Map Store’s maximal requirement to be less than 50K expert maps with 200 MB CPU memory (Xue et al., 2024).
4.5. Expert Caching and Eviction
Similar to existing expert offloading solutions (Eliseev and Mazur, 2023; Xue et al., 2024; Song et al., 2024), we design FineMoE to maintain an Expert Cache on GPUs to reuse expert weights when serving different request prompts. Given searched expert maps from §4.2, we guide FineMoE’s Expert Cache to compute two priority scores for individual experts: 1) a prefetching priority to decide the orders to prefetch experts in the searched maps, and 2) an eviction priority to determine the orders to evict experts in the Expert Cache.
@@ 4.5. Expert Caching and Eviction
where is the expert probability from the searched expert map, and is the current layer that the inference process stays at. Intuitively, experts with a higher probability to be activated should be prefetched sooner, and experts that sit closer to the current layer (i.e., smaller ) should also be prioritized.
@@
Expert eviction priority. Similar to MoE-Infinity (Xue et al., 2024), FineMoE’s expert caching is based on the least frequently used (LFU) caching algorithm. We integrate the searched map to jointly determine the eviction priority. For each expert , we define the eviction priority to be
@@
where is the cache visit frequency and is the probability from the searched map for an expert . Intuitively, when reaching the Expert Cache limit, we want to first evict experts who are less frequently hit and have lower probabilities of being activated. Note that similar to existing works (Xue et al., 2024; Song et al., 2024), we do not consider the recent usage of experts as opposed to the classic least recently used (LRU) algorithm (Eliseev and Mazur, 2023). Since the expert usage is layer-wise sequential, i.e., one layer following another, prioritizing recently used experts is against the nature of sequential forward computation.
@@
On-demand expert loading. Mispredictions of expert prefetching lead to expert miss in the Expert Cache, as the MoE model cannot find available experts designated by the gate networks. Whenever an expert miss occurs, FineMoE pauses all expert prefetching tasks and immediately loads missed experts from CPU to GPU memory for fast serving.
5. FineMoE’s Implementation
@@
We prototype FineMoE on top of Huggingface Transformers framework (Wolf et al., 2020) using MoE-Infinity codebase (Xue, Leyang and Fu, Yao and Lu, Zhan and Mai, Luo and Marina, Mahesh, [n. d.]). The implementation of FineMoE is described as follows.
@@
Expert Map Store is implemented in Python using PyTorch (Paszke et al., 2019) and NumPy (Harris et al., 2020) libraries. We store all semantic embeddings and expert maps using ndarrays data structure for efficient array operations. The arrays are converted to tensors to compute similarity for expert map searching.
Expert Map Searcher is implemented in Python using PyTorch (Paszke et al., 2019). We implement the pairwise computations, including similarity (§4.2) and redundancy (§ 4.4) scores, using PyTorch native operations.
@@
Expert Cache is implemented in C++ based on MoE-Infinity codebase (Xue, Leyang and Fu, Yao and Lu, Zhan and Mai, Luo and Marina, Mahesh, [n. d.]). The expert management in GPUs is implemented with the CUDA Runtime APIs (NVIDIA, 2024). We implement prefetching and caching logic of FineMoE in the MoE-Infinity codebase to enable expert offloading. Same with MoE-Infinity, FineMoE supports multi-GPU inference with expert parallelism (EP), where the experts are mapped to different GPU devices for loading and offloading. We use a hash map to assign expert IDs to different GPUs and retrieve them during inference. The expert assignment follows a round-robin manner to balance the overall GPU load. Additionally, we use a task pool in the GPU space with asynchronous threads to schedule and execute expert prefetching and on-demand loading tasks.
6. Evaluation
6.1. Experimental Setup
We introduce our evaluation methodology in this section.
@@ 6.1. Experimental Setup
Testbed. We conduct all experiments on a six-GPU testbed, where each GPU is an NVIDIA GeForce RTX 3090 with 24 GB GPU memory. All GPUs are interconnected using pairwise NVLinks and connected to the CPU memory using PCIe 4.0 with 32GB/s bandwidth. Additionally, the testbed has an AMD Ryzen Threadripper PRO 3955WX CPU with 32 cores and 480 GB CPU memory.
@@
Models. We employ three popular MoE-based LLMs in our evaluation: Mixtral-87B (Jiang et al., 2024b), Qwen1.5-MoE (Yang et al., 2024), and Phi-3.5-MoE (Abdin et al., 2024). Table 1 describes the parameters, number of MoE layers, and number of experts per layer for the three models. Following the evaluation of existing works (Song et al., 2024), we profile the models to set the optimal prefetch distance to three before evaluation.
@@
Datasets and traces. We employ two real-world prompt datasets commonly used for LLM evaluation: LMSYS-Chat-1M (Zheng et al., 2023) and ShareGPT (ShareGPT, 2022). For most experiments, we split the sampled datasets in a standard 7:3 ratio, where 70% of the prompts’ context data (i.e., semantic embeddings and expert maps) are stored in FineMoE’s Expert Map Store, and 30% of the prompts are used for testing. For online serving experiments, we empty the Expert Map Store and use real-world LLM inference traces (Patel et al., 2024; Stojkovic et al., 2025) released by Microsoft Azure to set input and generation lengths and drive invocations.
@@
Baselines. We compare FineMoE against four SOTA MoE serving baselines: 1) MoE-Infinity (Xue et al., 2024) uses coarse-grained request-level expert activation patterns and synchronous expert prediction and prefetching for MoE serving. We prepare the expert activation matrix collection for MoE-Infinity before evaluation for a fair comparison. 2) ProMoE (Song et al., 2024) employs a stride-based speculative expert prefetching approach for MoE serving. Since the codebase of ProMoE is not open-sourced and requires training predictors for each MoE model, we reproduced a prototype of ProMoE on top of MoE-Infinity in our best effort. 3) Mixtral-Offloading (Eliseev and Mazur, 2023) combines a layer-wise speculative expert prefetching and a LRU-based expert cache. 4) DeepSpeend-Inference (Aminabadi et al., 2022) employs an expert-agnostic layer-wise parameter offloading approach, which uses pure on-demand loading and does not support prefetching. We implement the offloading logic of DeepSpeed-Inference in the MoE-Infinity codebase and add an expert cache for a fair comparison. We enable all baselines to serve MoE models from HuggingFace Transformer (Wolf et al., 2020).
@@
Metrics. Following the standard evaluation methodology of existing works (Song et al., 2024; Xue et al., 2024; Zhong et al., 2024; Agrawal et al., 2024) on LLM serving, we report the performance of the prefill and decode stages separately. We measure Time-to-First-Token (TTFT) for the prefill stage and Time-Per-Output-Token (TPOT) for the decode stage. Additionally, we also report other system metrics, such as expert hit rate and overheads, for detailed evaluation.
6.2. Offline Serving Performance
We first evaluate the offline serving performance of prefill and decode stages when running FineMoE and other baselines with the three MoE models, where we report Time-To-First-Token (TTFT) and Time-Per-Output-Token (TPOT). Similar to existing works (Zhong et al., 2024; Agrawal et al., 2024), we measure TTFT and TPOT for individual prompts for each combination of model and dataset. For evaluation with LMSYS-Chat-1M and ShareGPT datasets, the input lengths are set to 37 and 43 tokens, and generation lengths to 127 and 122 tokens, which are the mean values calculated across datasets, respectively. For each dataset, we randomly sample 64 prompts and report average results.
@@ 6.2. Offline Serving Performance
Figure 10 shows the TTFT, TPOT, and expert hit rate of FineMoE and four baselines when serving three MoE models with LMSYS-Chat-1M and ShareGPT datasets, respectively. DeepSpeed-Inference has both the worst TTFT and TPOT due to expert-agnostic offloading and lacking expert prefetching. While Mixtral-Offloading, ProMoE, and MoE-Infinity perform better than DeepSpeed-Inference, they are underperformed by FineMoE because of coarse-grained offloading designs. Compared to DeepSpeed-Inference, Mixtral-Offloading, ProMoE, and MoE-Infinity, FineMoE reduces the average TTFT by 74%, 67%, 56%, and 53%, and reduces the average TPOT by 46%, 38%, 27%, and 22%, respectively.
For expert hit rate, DeepSpeed-Inference has no expert misses because it fetches whole layers with full experts, but with the worst latency due to pure on-demand loading. Mixtral-Offloading achieves a higher hit rate than ProMoE and MoE-Infinity because of its synchronous speculative prefetching with a prefetch distance of 1. However, due to synchronous prefetching, its TTFT and TPOT are worse than others except DeepSpeed-Inference. Overall, FineMoE improves the average expert hit rate by 14%, 37%, and 68% over Mixtral-Offloading, ProMoE, and MoE-Infinity, respectively.
6.3. Online Serving Performance
@@
Except for the offline evaluation (i.e., Expert Map Store in full capacity before serving), we also evaluate FineMoE against other baselines in online serving settings. We empty the Expert Map Store of FineMoE and the expert activation matrix collection of MoE-Infinity for the online serving experiment. The request traces are derived from Azure LLM inference traces (Patel et al., 2024; Stojkovic et al., 2025), with randomly sampled 256 requests (2.91 requests per second), to drive LMSYS-Chat-1M prompts for each MoE model serving. To ensure consistency, FineMoE and all baselines input and generate the exact number of tokens specified in the traces. Figure 11 illustrates the CDF of end-to-end request latency across three MoE models. The results demonstrate that FineMoE significantly reduces overall request latency compared to other baselines in online serving.
6.4. Impact of Expert Cache Limits
@@
We measure the TPOT of FineMoE and other baselines by limiting the expert cache memory budget to investigate their performance in the latency-memory trade-off (§2.3). We mainly focus on TPOT to show the end-to-end performance impacted by varying cache limits. Figure 12 shows the TPOT of FineMoE and four baselines when serving three MoE models under different expert cache limits. We gradually increase the GPU memory allocated for caching experts from 6 GB to 96 GB while employing the same experimental setting in §6.2. Similarly, DeepSpeed-Inference has the worst TPOT due to being expert-agnostic. FineMoE consistently outperforms Mixtral-Offloading, ProMoE, and MoE-Infinity under varying expert cache limits. As the cache limit increases, the performance gap between all baselines narrows due to the increased availability of cached experts. Nevertheless, for limited GPU memory sizes (e.g., 6GB), FineMoE reduces the TPOT by 36%, 25%, 16%, and 29%, compared to DeepSpeed-Inference, Mixtral-Offloading, ProMoE, and MoE-Infinity, across three MoE models, respectively. With fine-grained expert offloading, FineMoE significantly reduces the expert on-demand loading latency while maintaining a lower GPU memory footprint, therefore achieving a better spot in the latency-memory trade-off of MoE serving.
6.5. Impact of GPU Performance
To evaluate the impact of GPU performance on offloading methods, we repeat the experiments using LMSYS-Chat-1M on an NVIDIA A100 testbed equipped with 80 GB of HBM2e memory and a peak memory bandwidth of 2 TB/s. Figure 13 presents the serving performance of FineMoE and the baselines across the three MoE models. FineMoE achieves smaller performance gains on the A100 than on the 63090 testbed, since high-end GPUs and the lack of EP yield faster inference and lower offloading overhead. Nevertheless, FineMoE consistently outperforms all baselines. The expert hit rate remains largely unaffected, as GPU performance has less impact on expert predictions.
6.6. Ablation Study
We present the ablation study of FineMoE’s design.
@@ 6.6. Ablation Study
Effectiveness of expert map search. One of FineMoE’s key designs is the expert map, which tracks expert selection preferences in fine granularity. We evaluate the effectiveness of the expert map against five expert pattern-tracking approaches as follows. 1) Speculate: speculative prediction used by Mixtral-Offloading (Eliseev and Mazur, 2023) and ProMoE (Song et al., 2024), 2) Hit count: request-level expert hit count used by MoE-Infinity (Xue et al., 2024), 3) Map (T): expert map with only trajectory similarity search, 4) Map (T+S): expert map with both trajectory and semantic similarity search but statically select top-K experts to prefetch, and 5) Map (T+S+): expert map with full features enabled, including trajectory and semantic similarity search (§4.2) and dynamically selecting experts to prefetch (§4.3). We implement the above methods in FineMoE’s Expert Map Searcher for a fair comparison. Figure 14(a) shows the expert hit rate of the above expert pattern tracking methods. Speculative prediction is effective due to the widespread presence of residual connections in Transformer blocks. However, its effectiveness decreases drastically as prefetch distance increases (Song et al., 2024). The request-level expert activation count has the worst performance due to coarse granularity. As features are incrementally restored to FineMoE’s expert map, the expert hit rate gradually increases, demonstrating its effectiveness.
@@
Effectiveness of expert prefetching and caching. We evaluate FineMoE’s expert prefetching and caching against two caching algorithms: 1) LRU used by Mixtral-Offloading (Eliseev and Mazur, 2023) and 2) LFU used by MoE-Infinity (Xue et al., 2024). Figure 14(b) depicts the expert hit rate of FineMoE and two baselines. The results show that LRU performs poorly in expert offloading scenarios. Though LFU achieves a higher hit rate than LRU, FineMoE surpasses both, achieving the highest expert hit rate.
6.7. Sensitivity Analysis